OpenStudy (anonymous):

@Holly00d1248 @dtan5457 Which expression is equal to a + (b + c)? (a + b) + c (a + b) ⋅ c a + bc b + ac

OpenStudy (holly00d1248):

Id go with C im not 100% sure though

OpenStudy (anonymous):

Commutative or Associative property, do you know what they are?

OpenStudy (anonymous):

(a+b)+c = a+(b+c) = b+(a+c)